Assume that having a dimple in your chin is a recessive trait, and can be labeled dd. Find the probability that the offspring of dad Dd, and mom Dd will have a child who has a dimpled chin. Show your work.
Answer #1

P( offspring of dad Dd, and mom Dd will have a child who has a dimpled chin)=P(dd trait from parents)

=P(d trail from Father and d trait from mother )=P(d trait from mother)*P(d trail from father)=(1/2)*(1/2) =1/4